Write the equation for the following relation. D = {(x, y): (3, 9), (5, 25), (-2, 4), . . .}
Answer by jim_thompson5910(35256)

Notice how 3^2 = 9, 5^2 = 25 and (-2)^2 = 4

So the equation is y = x^2